OpenStudy (anonymous):

Given points A(3, 2) and B(9, 1). Answer the following question concerning using the Pythagorean Theorem to find AB. What is this value of the horizontal leg?

OpenStudy (anonymous):

horizontal leg?

OpenStudy (anonymous):

thats all they gave me

OpenStudy (anonymous):

ooohh ok i see

OpenStudy (anonymous):

horizontal leg is 1

OpenStudy (anonymous):

you are supposed to connect the dots and then draw a right triangle. the base will be 1 (the distance from 1 to 2 in the y) and the vertical will be 6, the distance from 3 to 9

OpenStudy (anonymous):

then you will be asked to find the length of the hypotenuse which is the distance between the two points. that will be \[\sqrt{1^2+6^2}=\sqrt{1+36}=\sqrt{37}\]